The demand curve for a good is Q= 1000-2p squared

What is the elasticity at the point

p=​$10.00 and Q=800​?

XText Question 1.5 The demand curve for a good is a-1,000-2p What is the elasticity at the point p $10.00 and Q 800? The elasticity of demand is ε-Π (Enter your response rounded to three decimal places and include a minus sign)
Answer #1


Demand curve is as follows -

Q = 1,000 - 2p2

Calculate the \Delta Q/\Delta P -

\Delta Q/\Delta P = dQ/dp = d(1,000 - 2p2)/dp = -4p

p = 10

So,

\Delta Q/\Delta P = -4p = -4 * 10 = -40

Calculate the elasticity of demand -

Elasticity of demand = (\Delta Q/\Delta p) \times \left ( p/Q \right )

Elasticity of demand = (-40) * (10/800) = -400/800 = -0.500

The elasticity of demand is -0.500